OpenStudy (anonymous):

the point (7,24) is on the terminal side of theta of an angle in standard position .determine the exact value of cot theta .

jimthompson5910 (jim_thompson5910):

hint: 

OpenStudy (anonymous):

its 7/24

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

but arent we suppose to find the hypotenuse side

jimthompson5910 (jim_thompson5910):

no need

OpenStudy (anonymous):

oh never mind i c

jimthompson5910 (jim_thompson5910):

cotangent doesn't use it, so it's just extra unused info

OpenStudy (anonymous):

k thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

question if u put tan 96 deg24 sec in your calculator do u get -9.1457 or -8.9152 because me and my friends got different answers .2 of us got the same answer and the other to got the other

jimthompson5910 (jim_thompson5910):

96 degrees, 24 seconds = 96 degrees + 0 minutes + 24 seconds 96 degrees, 24 seconds = 96 degrees + 0 minutes + 24/3600 degrees 96 degrees, 24 seconds =( 96 + 24/3600) degrees

OpenStudy (anonymous):

yeah but ther want us to put it in decimal notation..i missed that out

OpenStudy (anonymous):

sorry

jimthompson5910 (jim_thompson5910):

tan(96 degrees, 24 seconds ) = tan( 96 + 24/3600) = ???

OpenStudy (anonymous):

tan 96degrees24minutes .round off to the nearestt 4 decimal places

jimthompson5910 (jim_thompson5910):

oh minutes

jimthompson5910 (jim_thompson5910):

tan(96 + 24/60) = ???

OpenStudy (anonymous):

i got -9.1457

jimthompson5910 (jim_thompson5910):

tan(96 + 24/60) = -8.91520085005536 so not sure how you got that

OpenStudy (anonymous):

thanks my calculator doesnt give me that answer but thanks

jimthompson5910 (jim_thompson5910):

hmm odd

jimthompson5910 (jim_thompson5910):

you're in degree mode right?